A tractor seat is mounted on springs. When a 78 kg adult sits on the seat, the characteristic frequency is 3 Hz. What is the characteristic frequency when a child of mass 22 kg sits on the seat? (Neglect the seat mass.)
1
Expert's answer
2020-03-27T09:07:06-0400
"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}"

So,


"f'=f\sqrt{\frac{m}{m'}}=3\sqrt{\frac{78}{22}}=5.65\ Hz"